OpenStudy (anonymous):

Solve x2 + 4x − 4 = 0. x equals the quantity of negative 2 plus or minus square root 16 all over 2 x equals negative 4 plus or minus square root 2 x equals negative 2 plus or minus 2 square root 2 x equals the quantity of 4 plus or minus square root 18 all over 2

OpenStudy (anonymous):

okay so you have to isolate x do you know how to do that?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

@scenemunster its C

OpenStudy (anonymous):

Use the quadratic formula.\[\huge~\rm~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

yep ^^

OpenStudy (anonymous):

im stuck at \[-4 \pm \sqrt{\frac{ 32 }{ 2 }}\]

OpenStudy (anonymous):

ôkay welll to start of you have to divide the entire -4+/-srt(32) by 2 not just the 32 in the square root after that you simply simplify

OpenStudy (anonymous):

i got c

OpenStudy (anonymous):

Which of the following is a solution of x2 + 2x + 8? 4 + i times the square root of 7 −1 + i times the square root of 7 1 + i times the square root of 7 −4 + i times the square root of 7

OpenStudy (anonymous):

im stuck on this one i got to \[-2\pm \sqrt{\frac{ 36 }{ 2 }}\]

OpenStudy (anonymous):

You add 32 and 4. :D Continue.